Solve the system x + 2y = -13
and 4x - 3y = 25 using substitution
method. Please include complete solutions.

Answer by Alan3354(69443)

Solve the system x + 2y = -13
and 4x - 3y = 25 using substitution
method.
--------------
x + 2y = -13 ---> x = -2y - 13
--
4x - 3y = 25
Sub for x
4*(-2y-13) - 3y = 25
-8y - 52 - 3y = 25
-11y = 77
y = -7
====================
x + 2y = -13
x + 2*(-7) = -13
x -14 = -13
x = 1